# Comparison of two Cooperative Maneuver Planning Approaches at a Real-World T-Junction

### v2

@@ Abstract
Connected automated driving promises a significant improvement of traffic efficiency and safety on highways and in urban areas.
Cooperative maneuver planning may facilitate active guidance of connected automated vehicles at intersections.
Research in automatic intersection management put forth a large body of works that mostly employ rule-based or optimization-based approaches primarily in fully automated simulated environments.
In this work, we compare two cooperative planning approaches for unsignalized intersections that are capable of handling mixed traffic, i.e., the road being shared by automated vehicles and regular vehicles driven by humans.
The first approach is a cooperative planner that selects the most efficient out of multiple possible maneuvers based on a scene prediction trained on real driving data.
The second cooperative planning approach is based on graph-based reinforcement learning, which conquers the lack of ground truth data for cooperative maneuvers.
We thoroughly evaluate both cooperative planners in a realistic high-fidelity simulation with fully automated traffic and mixed traffic.
The simulative experiments show that cooperative maneuver planning leads to less delay due to interaction and a reduced number of stops.
Furthermore, we present results from real-world experiments with three prototype automated vehicles at a T-junction in public traffic, in which both planning modules demonstrate their ability to perform efficient cooperative maneuvers.
6.89in(1.81in,1.7in)
@@
Connected automated driving, cooperative planning, driver model, scenario prediction, optimization, reinforcement learning, graph neural network, mixed traffic, simulation, real-world evaluation, traffic efficiency.
1 INTRODUCTION
Urban traffic is prone to inefficiencies and disturbances due to the ever increasing volume of traffic. This manifests, e.g., at smaller intersections, where static priority rules are the prevailing method for coordination of vehicles. Connected automated driving opens up new opportunities to improve urban traffic efficiency by leveraging communication links between vehicles and possibly infrastructure systems. Moreover, edge computing resources are becoming available in urban areas that enable, e.g., populating and maintaining of a collective environment model (EM) of an area of interest. This setup allows connected automated vehicles (CAVs) to improve their local planning algorithms by incorporating data from the server-side EM, as it was shown in [1] at a suburban three-way intersection in Ulm-Lehr, Germany.
Connectivity between vehicles can also be used for the coordination of maneuvers to increase traffic efficiency and safety. In the current work, we integrate and evaluate two cooperative crossing maneuver planners, as illustrated in Fig. 1. The cooperative maneuvers are planned on a centralized edge server, which instructs CAVs to explicitly deviate from priority rules to reach peak efficiency. As urban automotive traffic will not be fully automated in the near future, mixed traffic, i.e., the simultaneous use of roads by both human-driven vehicles (HDVs) and CAVs, will be prevalent. Therefore, any real-world cooperative planner must consider HDVs that cannot be directly influenced and behave according to the priority rules. We compare two approaches to cooperative maneuver planning, one of which leverages a multi-scenario prediction to derive an optimal maneuver, based on [2]. The second approach employs reinforcement learning (RL) to train a graph neural network (GNN) policy for cooperative maneuver planning [3]. To provide realistic results, we employ the real-time simulation framework DeepSIL [4] with a state-of-the-art human driver model for HDVs. We extended the simulation framework to support multiple CAVs and integrated both maneuver planners as well as a trajectory planner and controller combination [5] for CAVs, the latter of which is also used in the real-world experiments. The execution of the cooperative maneuvers in simulation and in real-world is based on the coordination protocol proposed in [6]. The core contribution of the present work is twofold:
Statistical analysis and evaluation of both planning approaches in a holistic simulation framework for fully automated and mixed traffic with realistic HDV and CAV models;
@@ 1 INTRODUCTION
The remainder of the paper is structured as follows: Section 2 discusses the state of the art in classical and learning-based cooperative planning for automated driving. The optimization-based maneuver planner is presented in Sec. 4, followed by the RL-based one in Sec. 5. Afterwards, we discuss our simulative evaluation results (Sec. 6) and real-world experiments in public traffic (Sec. 7). Finally, Sec. 8 summarizes the results of the article and gives an outlook on future work.
2 RELATED WORK
@@
The enumeration of all possible vehicle crossing orders in CAV-only traffic yields to the optimal ordering [8]. However, the number of order combinations is exponential in the number of vehicles as shown in [9], thus, exhaustive methods are not feasible in a real-time system. A more efficient approach is to use a heuristics to directly generate a single crossing order. One of the earliest proposed heuristics is a first-come, first-served scheme to reserve cells in a discretized space-time grid for each vehicle, which was used by several research groups [10, 11]. Other approaches try to find a near-optimal crossing order of the arriving vehicles using more elaborated algorithms. Possible optimizing strategies are dynamic programming [12], ant colonies [9], a control policy based on Petri Nets [13], Mixed Integer Quadratic Programs [14], or Monte Carlo Tree Search [15]. However, these works are designed for CAV-only traffic, and only the authors in [13] show the real-time capability of their algorithm in simulation and in real-world experiments. The other methods are tested in synthetic non-real-time simulations only, not considering realistic traffic behavior and processing delays.
Fewer researchers have tackled the intersection management task for much more challenging mixed traffic and especially scenarios with a vast majority of HDVs, which will be prevalent in the near future. Previous mixed traffic approaches employ signalized intersection management to support the presence of HDVs. Dresner and Stone extended their cell reservation system from previous works by different fallback modes in [16]. Those modes use dynamically controlled traffic light phases to allow safe crossing of a low rate of HDVs (). Bento et al. employ a similar system that switches traffic lights to green for individual vehicles [17], which works well for a low percentage of HDVs. For both approaches, the gains deteriorate in an environment with a lot of HDVs. Two different research groups [18, 19] proposed virtual platoons for signalized intersection management supporting mixed traffic at high HDV rates. Their methods use a heuristics and a branch-and-bound algorithm, respectively, and are evaluated in a traffic simulation demonstrating real-time capability. However, all these methods require traffic lights to control HDVs. Also, most approaches assume that the turn direction of HDVs is known, which is generally not given in real-world traffic.
The approach proposed in our previous work in [2] is specifically designed to work at unsignalized intersections with any HDV rate and unknown HDV turn directions. It employs a scene-consistent multi-scenario prediction, which we enhanced for this paper, and optimizes the expected efficiency by maneuvers between the present CAVs. A respective coordination protocol to execute the cooperative maneuvers on CAVs was previously proposed in [6], making this approach applicable on real vehicles.
@@
In [22], the authors propose an RL approach assuming that automated vehicles only perceive their immediate surroundings. The RL policy is shown to manage intersection traversals in presence of simulated human-driven vehicles while turning maneuvers are disallowed. A further RL framework that combines local observations with a joint reward to accommodate a cooperative objective was presented by [23]. Although this model is capable of handling mixed traffic, the network architecture limits the maximum number of cooperating vehicles in the scene. These works also lack the application to real vehicles and a comprehensive evaluation in public traffic.
@@
3 COOPERATIVE MANEUVER PLANNING MODULES
In our proposed system architecture, the centralized cooperative planning module on the edge server plans a joint maneuver on behavior level, which is distributed to all CAVs in the scene. Each of the CAVs runs a local motion planner to derive a viable trajectory. Typical motion planning algorithms require a specification of the desired motion for a sufficient planning horizon (at least multiple seconds) into the future. Therefore, the distributed maneuver representation shall enable the vehicles to individually plan their trajectory locally.
Basis for the planning of cooperative maneuvers is the server-side EM. This EM is the result of fusing multiple data sources, like cooperative awareness messages (CAM, [27]) and collective perception messages (CPM, [28]) from connected vehicles or infrastructure perception [1]. Thus, the server-side EM is more extensive than the individual CAVs’ ones and is considered to contain all information relevant for maneuver planning. A vehicle in the EM is denoted as
@@ 3 COOPERATIVE MANEUVER PLANNING MODULES
The derived maneuvers are passed to the CAVs according to the maneuver coordination protocol proposal in [6]. While the protocol is very generic and supports multiple cooperative use cases, we focus on the components that are relevant for AIM. Thus, the final step to provide the planned cooperative maneuver to the CAVs is to derive so-called longitudinal maneuver waypoints defined as
where its position is given by . The interval denotes the admitted time window for the vehicle to cross the waypoint. The fields and describe the preceding and following vehicle IDs, respectively. Note that lateral guidance is provided by lane centerlines in a common map that is available to all vehicle-side motion planners. This additional information can be leveraged by vehicle-side motion planners to improve follow trajectory planning on lead vehicles.
4 OPTIMIZATION-BASED PLANNER
The optimization-based planner builds upon our previous works presented in [2]. The main component is our scene-consistent prediction module [29] that estimates the behavior and velocities of the vehicles in the currently observed traffic scene in the immediate future. Multiple such scenario predictions with various applicable maneuvers are performed and the best one—according to validity and efficiency metrics—is selected to be communicated to the connected vehicles.
The planning algorithm uses the concept of conflicting vehicles and conflict zones. Two vehicles are in conflict when they are on merging or crossing lanes and both still approach this critical point. Conflict zones (cf. Fig. 2) are areas where two conflicting vehicles might collide, therefore only vehicles from one direction may occupy a conflict zone at any time. The conflict zones are extracted geometrically from the map data.
Connected vehicles share their routes with the planning module, therefore their future turn direction is known to the algorithm. HDVs, however, only use their turn indicators to communicate their turn direction, which cannot be reliably detected and used in the planning module. Therefore, it is conservatively assumed that an HDV may be in conflict with any crossing or merging lanes, and the CAVs in the underlying prediction need to yield respectively. Vice-versa, we assume that HDVs do not know the route of other vehicles or do not rely on their turn indicators, thus all predicted HDVs expect a conflict with any crossing or merging lanes as well.
Additionally, it is assumed for the prediction that CAVs will accept a proposed maneuver. Whenever a request is rejected, the current maneuver is aborted for all vehicles and the planning state is reset.
4.1 Driver Model for Prediction
Vital to the maneuver planning is the scenario prediction module. We employ a longitudinal lane-based prediction module trained on real-world traffic data. The driver model comprises two multi-layer perceptrons (MLPs): estimates the acceleration during the next timestep whereas determines the gap acceptance, i.e., the decision whether a vehicle should enter the intersection before another prioritized vehicle. The driver model is evaluated jointly for all vehicles in the traffic scene and integrated over a horizon with a time step of for each evaluated scenario.
@@ 4.1 Driver Model for Prediction
Current and maximum velocity | , |
@@
Distance to point of guaranteed arrival | |
Velocity | |
Distance of other vehicle to stop line | |
Velocity of other vehicle |
and are the observations of towards its environment and towards a conflicting vehicle , respectively. The values are shown in Table 1 and Fig. 3.
The acceleration estimation has 2 hidden layers of size 16 and a activation function. The observations and the gap acceptance form the inputs of the network:
The model was trained using proximal policy optimization (PPO) [30] in an RL closed-loop simulation environment. The used objective function encourages driving near the speed limit and penalizes collisions with prioritized vehicles to emulate natural driving behavior.
The gap acceptance model has 2 hidden layers of size 16 and a LeakyReLU activation function. The decision of whether to enter the intersection is given by the minimum of the gap acceptance against every conflicting vehicle , where assigned priorities can override the MLP output:
The model was trained in a supervised learning fashion on real driving data extracted from the inD dataset [31].
4.2 Priority Assignment Search
The internal maneuver representation is a list of priority assignments between pairs of cooperative vehicles. Each entry is a pair of vehicles , where the first is prioritized over the second. The planner operates cyclically in a frequency of and tries to add assignments to the resulted from previous cycles to optimize the efficiency.
@@ 4.2 Priority Assignment Search
where denotes the lane speed limit of vehicle at its position at time .
At the beginning of a planning cycle , the current base efficiency is determined by predicting the current scene with the priority assignments from the last cycle, where vehicles that already left the intersection are removed. Next, the current scene is predicted another time without any assigned priorities. If this leads to a lower time loss, the previous maneuver is not efficient anymore, due to some of the vehicles driving differently than predicted. In this case, the maneuver is terminated () and the planning state is reset.
Otherwise, it is attempted to extend by one additional priority assignment. For each conflicting pair of CAVs, the efficiency of both and is evaluated. This means that for conflicting CAV pairs an additional number of scenarios needs to be predicted. For (which is nearly always reached at a typical intersection), simultaneously extending by two or more additional priority assignments is also evaluated. The total number of predicted scenarios is limited to to retain a real-time capable planning scheme.
@@
Each scenario prediction is checked for freedom from collision and for correct crossing order according to the respective priority assignments. The most efficient of the valid priority assignment lists is selected as the result of the current planning cycle.
4.3 Derivation of the Cooperative Maneuver
The priority assignments are descriptive enough from a centralized point of view, but need to be converted to maneuver constraints as in Eq. (2) for each individual CAV. Each pair of conflicting CAVs shares a conflict zone, i.e., a segment along the road that may only be occupied by one of the vehicles at any time to guarantee freedom from collision, as explained above. The predicted timestamp at which the prioritized vehicle leaves and the yielding vehicle may enter the conflict zone is used to encode the space-time maneuver constraints. The start and end points of the conflict zone along the respective vehicle routes are used as waypoints for the constraints.
5 LEARNING-BASED PLANNER
This section introduces our proposed learning model (Sec. 5.1) for cooperative planning in mixed traffic, the graph-based representation (Sec. 5.2), and the reward function (Sec. 5.3). Finally, Sec. 5.4 discusses the integration of the RL policy with dedicated motion planning algorithms. We combine the methods presented in our previous works [3] and [26] and provide a brief summary of the main features.
5.1 Learning Model
Due to the lack of ground-truth data for cooperative maneuvers in urban traffic, we consider the planning task a multi-agent RL problem. In a multi-agent RL setting, various learning paradigms exist depending on the degree of centralization [32]. Joint cooperative planning in connected automated driving is best modeled by the centralized training centralized execution (CTCE) paradigm. Thus, we define the multi-agent planning task as a single partially observable Markov decision process (POMDP), denoted as
The traffic environment in the area of operation is denoted by a state . This includes the full state information on all CAVs and HDVs. The environment is transferred into the next state by applying an action from the joint action set. denotes the transition function that yields the probability of changing from to under action as . Because human reaction to a given traffic state is not deterministic, the behavior of HDVs is neither, which manifests in a nondeterministic state transition. The desired policy behavior is provided through a scalar reward signal given by the function , whose detailed setup is discussed in Sec. 5.3. The cooperative planner cannot observe the entire abstract traffic state because, e.g., the maneuver intention of human drivers is unknown. A smaller observation set is defined to resemble the state information that is available to the cooperative planner through the communication link or infrastructure perception. The mapping is not injective and additionally models the influence of measurement uncertainties.
@@ 5.1 Learning Model
Since the joint planning task is being modeled by a single POMDP, the dimensionality of the observation space depends on the number of vehicles currently in the scene and may vary over time. Moreover, the dimensionality of the action space depends on the number of cooperative vehicles present and requires a permutation equivariant mapping to the observation space. Therefore, the input representation has to be suited to encode the varying number of dynamically interacting agents efficiently.
5.2 Graph-based Representation and Network

### v5

# Real-World Evaluation of two Cooperative Intersection Management Approaches

@@ Abstract
Cooperative maneuver planning promises to significantly improve traffic efficiency at unsignalized intersections by leveraging connected automated vehicles. Previous works on this topic have been mostly developed for completely automated traffic in a simple simulated environment. In contrast, our previously introduced planning approaches are specifically designed to handle real-world mixed traffic. The two methods are based on multi-scenario prediction and graph-based reinforcement learning, respectively. This is the first study to perform evaluations in a novel mixed traffic simulation framework as well as real-world drives with prototype connected automated vehicles in public traffic. The simulation features the same connected automated driving software stack as deployed on one of the automated vehicles. Our quantitative evaluations show that cooperative maneuver planning achieves a substantial reduction in crossing times and the number of stops. In a realistic environment with few automated vehicles, there are noticeable efficiency gains with only slightly increasing criticality metrics.
Index Terms:
Connected automated driving, cooperative planning, driver model, scenario prediction, optimization, reinforcement learning, graph neural network, mixed traffic, simulation, real-world evaluation, traffic efficiency.white{textblock*}7.1in(0.7in,0.15in)
@@
I Introduction
Urban traffic is prone to inefficiencies and disturbances due to the ever-increasing volume of traffic. This manifests, e.g., at smaller intersections, where static priority rules are the prevailing method for coordination of vehicles. Connected automated driving opens up new opportunities to improve urban traffic efficiency by leveraging communication links between vehicles and possibly infrastructure systems. Moreover, edge computing resources are becoming available in urban areas that enable, e.g., populating and maintaining of a collective environment model (EM) of an area of interest. This setup allows connected automated vehicles (CAVs) to improve their local planning algorithms by incorporating data from the server-side EM, as it was shown in [4] at a suburban three-way intersection in Ulm-Lehr, Germany.
Connectivity between vehicles can also be used for the coordination of maneuvers to increase traffic efficiency and safety. In the current work, we integrate and evaluate two cooperative crossing maneuver planners, as illustrated in Fig. 1. The cooperative maneuvers are planned on a centralized edge server, which instructs CAVs to explicitly deviate from priority rules to reach peak efficiency. As urban automotive traffic will not be fully automated in the near future, mixed traffic, i.e., the simultaneous use of roads by both human-driven vehicles (HDVs) and CAVs, will be prevalent. Therefore, any real-world cooperative planner must consider HDVs that cannot be directly influenced and behave according to the priority rules. We compare two approaches to cooperative maneuver planning, one of which leverages a multi-scenario prediction to derive an optimal maneuver, based on [25]. The second approach employs reinforcement learning (RL) to train a graph neural network (GNN) policy for cooperative maneuver planning [19].
Our methods are designed for urban intersections with one lane per direction. Larger intersections are beyond the scope of this paper, as they are typically managed by traffic lights and require a different cooperative planning paradigm.
To achieve realistic and comparable quantitative results, we employ the real-time simulation framework DeepSIL [36] with a state-of-the-art human driver model for HDVs. We extended the simulation framework to support multiple CAVs, maneuver planning modules, and V2X communication in between. Also, we integrated the trajectory planner and controller combination from [32] as well as an advanced vehicle model from one of our real-world CAVs into the simulation. Additionally, to the best of our knowledge, we are the first to provide an experimental evaluation of cooperative planners in public mixed traffic using real test vehicles and V2X communication. The execution of the cooperative maneuvers in simulation and in real-world is based on the coordination protocol proposed in [26]. Thus, the core contribution of the present work is threefold:
Extension of our simulation framework to integrate cooperative maneuvers and realistic HDV and CAV models;
Statistical analysis and evaluation of our two planning approaches in fully automated and in mixed traffic;
@@ I Introduction
The remainder of the paper is structured as follows: Section II discusses the state of the art in classical and learning-based cooperative planning for automated driving. Our previously introduced maneuver planning approaches are summarized and compared in Section III. We give an in-depth overview over our simulation framework and evaluation method in Section IV. Afterwards, we discuss our simulative evaluation results (Section V) and real-world experiments in public traffic (Section VI). Finally, Section VII summarizes the results of the article and gives an outlook on future work.
II Related Work
@@
II-A Classical Cooperative Maneuver Planning
The earliest cooperative planning methods rely on an exhaustive vehicle ordering search [23] or a simple heuristic such as first-come, first-served [6, 41]. Later approaches try to find a near-optimal crossing order of the arriving vehicles using more sophisticated algorithms like dynamic programming [42], ant colonies [39], a control policy based on Petri Nets [1], Mixed Integer Quadratic Programs [15], or Monte Carlo Tree Search [21]. However, all of these works are designed for CAV-only traffic and most were evaluated in simplistic simulation environments. Only the authors in [1] show the real-time capability of their algorithm in simulation and in real-world experiments.
Fewer researchers have tackled the intersection management task in the much more challenging mixed traffic setting and especially scenarios with a vast majority of HDVs, which will be prevalent in the near future. To the best of our knowledge, all previous mixed traffic approaches need to incorporate traffic lights to support the presence of HDVs. The first notable approaches [7, 2] were designed for a low rate of HDVs ( 10 %) and the gains deteriorate in an environment with a higher amount of HDVs. Two later publications [29, 43] proposed virtual platoons for signalized intersection management, supporting mixed traffic at high HDV rates and being evaluated in a real-time traffic simulation. However, all methods mentioned above require traffic lights to control HDVs and as such are out of our problem scope. Also, most approaches assume that the turn direction of HDVs is known, which is generally not given in real-world traffic.
Our optimization-based approach proposed in [25] is specifically designed to handle unsignalized intersections with any HDV rate and unknown HDV turn directions. It employs a scene-consistent multi-scenario prediction and generates maneuvers between the present CAVs to optimize the expected efficiency. A respective coordination protocol to execute the cooperative maneuvers on CAVs was previously proposed in [26], making this approach applicable on real vehicles.
II-B Learning-based Cooperative Maneuver Planning
@@
In [30], the authors propose an RL approach managing intersection traversals in presence of simulated human-driven vehicles while turning maneuvers are disallowed. A further RL framework that combines local observations with a joint reward to accommodate a cooperative objective was presented by [12] and is capable of handling mixed traffic, though with a limited number of cooperating vehicles in the scene. Both works lack the application to real vehicles and a comprehensive evaluation.
@@
III Cooperative Maneuver Planning Modules
In this section, the commonalities and differences of the two evaluated maneuver planning modules are presented. In our proposed system architecture, all CAVs in the scene provide their current state and desired route in the form of a sparsely sampled polyline to the centralized cooperative planning module on the edge server. The planning module generates a joint maneuver on behavior level, which is in turn distributed to all CAVs.
III-A Definition and Coordination of Cooperative Maneuvers
For both planning modules, the maneuvers are passed to the CAVs according to the maneuver coordination protocol proposal in [26]. Depending on the number of vehicles on conflicting paths, up to two maneuver waypoints (one at the intersection entry and one at the intersection exit) are passed to a given CAV. Each vehicle remains responsible for its own motion planning but needs to fulfill constraints defined by the maneuver waypoint as
The interval denotes the admitted time window for the vehicle to cross the waypoint . The fields and describe the preceding and following vehicle IDs, respectively. This additional information can be leveraged by vehicle-side motion planners to improve follow trajectory planning on lead vehicles [32]. Note that lateral guidance is provided by lane centerlines in a common map that is available to all vehicle-side motion planners.
III-B Environment Model Processing
The basis for both cooperative maneuver planning modules is the server-side EM. This EM is the result of fusing multiple data sources, like cooperative awareness messages (CAM, [8]) and collective perception messages (CPM, [9]) from connected vehicles or infrastructure perception [4] and is considered to contain all information relevant for maneuver planning. A vehicle in the EM is denoted as
@@ III-B Environment Model Processing
In a preprocessing step, both planning modules associate and project the vehicles in the server-side EM to the lane centerlines of an internal map format. While connected vehicles share their routes with the planning module, the turn intention of human drivers is unknown. Thus, both planners conservatively assume that an HDV may be in conflict with any crossing or merging lanes. Additionally, it is assumed that HDVs do not know the route of other vehicles and that the indicator light states cannot be reliably evaluated. The handling of this uncertainty might be improved by employing a prediction algorithm such as [35], which is out of scope for this work, though. Moreover, vulnerable road users that are not lane-bound are not accounted for in the present setup.
III-C Optimization-based Planning Algorithm
The optimization-based planner builds upon our previous works presented in [25]. The main component is a scene-consistent prediction module that estimates the behavior of the vehicles of the current traffic scene in the immediate future. Multiple such scenario predictions with various applicable maneuvers are performed and the best one—according to the validity and efficiency metrics defined below—is selected.
III-C1 Driver Model for Prediction
Vital to the maneuver planning is the scenario prediction module that we previously published in [27]. We employ a longitudinal lane-based prediction module trained on real-world traffic data. The driver model comprises two multi-layer perceptrons (MLPs) with 2 hidden layers of size 16 each: estimates the acceleration during the next time step. determines the gap acceptance, i.e., the decision whether a side road vehicle should merge or cross in a gap in main road traffic. The driver model is evaluated jointly for all vehicles in the traffic scene and integrated over a 15 s horizon with a time step of 0.1 s. This prediction is run for each priority assignment as described in Section III-C2. The input features are shown in Table I and Fig. 2.
@@ III-C1 Driver Model for Prediction
| Current velocity and speed limit | , |
@@
| Distance to end of conflict zone | |
| Velocity of this and the other vehicle | , |
| Distance of other vehicle to stop line |
The acceleration estimation was trained using proximal policy optimization (PPO) [34] in an RL closed-loop simulation environment on different intersection layouts. The objective function encourages driving near the speed limit and penalizes collisions with prioritized vehicles to emulate natural driving behavior. The gap acceptance model was trained using supervised learning on real driving data extracted from the inD dataset [3].
III-C2 Priority Assignment Search
Each possible maneuver is internally represented as a list of priority assignments between pairs of cooperative vehicles. Each entry is a pair of vehicles , where the first is prioritized over the second. The planner operates cyclically in a frequency of 5 Hz and tries to add assignments to the previous cycle result to optimize the efficiency, by predicting the scenario for each .
@@ III-C2 Priority Assignment Search
where denotes the lane speed limit of vehicle at its position at time . The scenario prediction for each is checked for freedom from collision and for correct crossing order according to the respective priority assignments. The most efficient of the valid priority assignment lists is selected as the result of the current planning cycle.
III-D RL-based Planning Algorithm
The RL-based cooperative planner combines the methods presented in our previous works [19] and [20].
III-D1 Learning Model
@@
Due to the lack of ground-truth data for cooperative maneuvers in urban traffic, we consider the planning task a multi-agent RL problem. The joint planning task is modeled by a single partially observable Markov decision process. Thus, the dimensionality of the observation space depends on the number of vehicles currently in the scene and may vary over time. Moreover, the dimensionality of the action space depends on the number of cooperative vehicles present and requires a permutation equivariant mapping to the observation space.
The RL policy is trained in a simulated traffic environment, based on the open-source simulator Highway-env [22], which has been extended for multi-agent planning. It shall derive a joint action for all CAVs in the scene composed of longitudinal acceleration commands in continuous space. We employ the TD3 [11] actor-critic RL algorithm to train a GNN for the cooperative planning task.
III-D2 Graph-based Representation and Network